AN UNUSUAL INTERNAL ROTATION: THE THREE-MINIMA MOTION OF AR IN TRIFLUOROMETHANE-AR

\underline{W. CAMINATI, A. MARIS, S.MELANDRI}, P. G. FAVERO, P. OTTAVIANI, Dipartimento di Chimica ”G.Ciamician”, Universita’ di Bologna, Via Selmi I-40126, Italy.

The rotational spectrum of Trifluoromethane-Ar (CHF_{3}-Ar, see Fig. on the right) has been observed by free jet absorption millimeter wave spectroscopy. The near prolate top spectrum of the adduct suggests that: i) the location of Ar is out of the C_{3v} axis of CHF_{3}; ii) the barrier to internal rotation of Ar is not negligible. The rotational lines are split in equally intense A E component lines for the low K_{−1} transitions, and in triplets lines for the high K_{−1}, transitions (doubly overlapped because of the near prolate degeneracy) with intensities 2:1:1 for A:E^{+}:E^{−} components. Such a behaviour is in accord with Hersbach theory for internal rotation.^{a}

aD. R. Herschbach and J. D. Swalen J. Chem. Phys. 29, 761 (1958).